OpenStudy (anonymous):

i need help thanks Use the Laws of Exponents to simplify the expression four to the power of forty over four to the power of twenty Explain how you found your answer and write the answer as an exponent.

OpenStudy (anonymous):

432 it's to hard to explain

OpenStudy (anonymous):

\[\frac{4^{40}}{4^{20}}=4^{40-20}=4^{20}\] \[Using \frac{x^{a}}{x^{b}}=x^{a-b}\]